Canadian Dollar: Spot and Forward (C$/$)

Euro: Spot and forward ($/€)

Mid rates

Bid

Ask

Mid rates

Bid

Ask

spot

1.2645

1.2639

1.2651

1.2390

1.2387

1.2393

Forward

3-week

23

27

19

15

3-month

135

128

155

146

7a. How much $100 will cost you in Canadian Dollar and Euro? (Hint: you are buying USD).

7b. What are 3-week and 3-month forward bid and ask rate for Canadian Dollar and Euro?

7c. Based on information above, what are the profit margin for the traders between Canadian Dollar and Euro? (Hint: the bid-offer spread % is the profit margin for dealers; you need to find the cross bid-offer rate between Canadian Dollar and Euro; use American Terms to show cross-rate)
Answer #1

Answer 7 a) Buying USD at ASK price

$ 100 = 1.2651*100 C$ = 126.51 C$

$ 100 = 1.2393*100 C$ = 123.93 E

Answer 7 b)

Canadian Dollar: Spot and Forward (C$/$) Euro: Spot and forward ($/€)
Mid rates Bid Ask Mid rates Bid Ask
spot 1.265 1.2639 1.2651 1.239 1.2387 1.2393
Forward
3-week 1.2639+0.0023=1.2662 1.2651+0.0027=1.2678 1.2387+0.0019=1.2406 1.2393+0.0015=1.2408
3-month 1.2639+0.0135=1.2774 1.2651+0.0128=1.2779 1.2387+0.0155=1.2542 1.2393+0.0146=1.2539

Answer 7c) Exchange rate S(C$/€) = S(C$/$)*S($/€)

S(C$/€)Bid = 1.2639*1.2387 = 1.5656 C$/€

S(C$/€)ASK = 1.2651*1.2393= 1.5679 C$/€

bid-ask spread % = (ask-bid)/bid *100 = (1.5679-1.5656)/1.5656 *100 = 0.1469%